OpenStudy (anonymous):

how do you graph 5X - 2Y = -6

OpenStudy (anonymous):

5x+6=2y \[y=\frac{ 5x+6 }{ 2 }\] then give x=1 and find Y give x=0 and also find y give x=-1 and find y draw it you have the value of x and y three times you can draw it

OpenStudy (anonymous):

get it ??

OpenStudy (anonymous):

No l/

OpenStudy (anonymous):

no ;/ can you draw it out

OpenStudy (anonymous):

 sorry it's too bad :)))

OpenStudy (anonymous):

now get it??

OpenStudy (anonymous):

here you are :DDD

OpenStudy (anonymous):

5x - 2y = -6 to find the x intercept, sub in 0 for y 5x - 2(0) = -6 5x = -6 x = -6/5 (if it makes it easier to graph, -6/5 is the same as -1 1/5) to find the y intercept and the slope, put the equation in y = mx + b form where m is your slope and b is your y intercept. 5x - 2y = -6 5x + 6 = 2y -- divide by 2 5/2x + 6/2 = y 5/2x + 3 = y (rearrange it) y = 5/2x + 3 so the y intercept is 3 your slope is 5/2 now plot your intercepts (-6/5,0) and (0,3) start at the x intercept (-6/5,0) and since the slope is 5/2, go up 5 spaces and to the right 2 spaces, then up 5 and to the right 2. and you should cross the y intercept at (0,3)